Find a particular solution to the differential equation using the Method of Undetermined Coefficients. x(t) – 10x(t) + 25x
Answer #1

. ر - 0 - 2 = ھی من که تو » کم عمر 97/ - z کم سے Comfort ax+bx+ ca = pertedt 1-5 put) = 126² Chorrecterictic eqn de-10d +25ا ( جو م ہے ہم = ج م + 20 - :: ی کی کمی میں لمبی ہوتی ہے یا پر ریمہ کا ری میک کیجیے کہ ہر مرد کی مکمم م ( مرا به مرگ برای م